# Data Distributional Properties As Inductive Bias for Systematic Generalization

### v2

@@ 7 Mutual Information versus Diversity
Figure 8: Normalized Mutual Information in the Training Set vs Out-of-Distribution Accuracy for the shape task for datasets with 8, 64 and 216 colors. For lower NMI, there is greater out-of-distribution performance. However, for similar levels of NMI but differing levels of diversity, we observe that greater diversity increases performance. This suggests that the effect of diversity is not only related to changing the NMI in the dataset.
Previous work [1] has shown that there is a relation between the Normalized Mutual Information (NMI) of latent attributes and systematic generalization. Mutual information quantifies the amount of information one variable (e.g., shape) provides about another (e.g., color). Then, this value is normalized with the entropy of the distributions to produce the NMI. The closer NMI is to zero, the more independent these factors are. In other words, a near-zero mutual information score indicates that the dataset offers strong evidence to the model that shape and color should be treated as independent attributes.
Changing diversity in a dataset impacts its NMI, so we proceed to study this effect. We seek to decouple the effect of increased diversity from a dataset’s NMI. To modulate NMI for a given level of diversity, we produce datasets that assign colors that both cubes and cylinders see during training (common colors) and a set of colors that are only seen either by cubes or by cylinders (exclusive colors). Modulating the ratio of common colors within the dataset directly alters its NMI, while keeping the total number of colors constant.
To measure the NMI, we follow previous work [1] and focus on the NMI between the color and shape of all objects of the training dataset. It shows a Pearson correlation of -0.79. Figure 8 plots results for datasets with differing levels of diversity. For lower NMI, there is greater out-of-distribution performance. However, for similar levels of NMI but differing levels of diversity, we observe that greater diversity increases performance. This suggests that the effect of diversity is not only related to changing the NMI in the dataset but rather other properties. Finally, we study the geometry of representations induced by diversity.
@@ 8.1 Disentanglement
A desirable property for SG is disentangled representations [27, 28, 17]. Using DCI metrics [19], we focus on measures of Disentanglement and Completeness. Disentanglement measures how many latent attributes each dimension of the representation can predict. A score of 0 implies that each dimension has an equal predictive power for all latent attributes, while a score of 1 suggests that each dimension exclusively predicts a single latent attribute. Completeness measures the extent to which different dimensions of the representation predict each latent attribute. A score of 0 means each latent attribute is predicted equally by all dimensions, whereas a score of 1 indicates that only one dimension is responsible for predicting each latent attribute.
Figure 9 shows results for these metrics for representations for the color and shape task derived from models with different levels of color diversity. Increased diversity for shape task representations leads to both greater disentanglement and completeness, while the converse is true for color task representations.
@@
Figure 9: Disentanglement and Completeness for models trained on different levels of color diversity for vectors predicting the shape and color tasks. Both disentanglement and completeness increase for the shape task vectors, while the contrary happens for color task vectors.
We also complement our analysis by drawing from neuroscience-inspired studies that have demonstrated the potential of parallelism in representations for enhancing generalization [8, 30].
@@ 8.2 Parallelism in Representations
For instance, think on a classic example for analogy in DNNs: $king-man+woman=queen$ which implies: $king-man=queen-woman=royalness$.
@@
In this case, subtracting the abstract concept of gender gives us vectors that move in the same direction (an abstract property one might consider royalness) for two otherwise different vectors. Encoding this property consistently in a single direction allows the model to generalize that concept by way of analogy. Analogy is a powerful inductive bias for solving new tasks [47].
@@
We proceed to apply this same principle to the latent factors contained in our dataset. From a starting set of 1.024 images, we sample 3.500 pairs of representations from objects that share a common latent attribute. To represent each object we use the representation vector associated to the attribute we wish to study, which in this case will be either shape or color. We proceed to find for each of these pairs another pair of representations which represents a change in this attribute (i.e. for shape, from a pair of red and green cubes, we look for a pair of red and green cylinders). We substract these pairs and from the resulting pair we obtain a single value called a p-score [30]. This score is calculated by measuring the cosine of the angle between vectors in the resulting pair. We report the mean of this value over the 3.500 pairs averaged over 5 runs.
@@
Figure 10: Computation of the p-score metric for vectors $V_{A}$ and $V_{B}$, representing the effect of ablating features on shape embeddings. We first subtract the embeddings to obtain vectors $V_{A}$ and $V_{B}$ and then calculate the degree of parallelism between them using cosine distance. A high p-score indicates strong parallelism, suggesting that the attribute (in this case, cubeness) is encoded consistently in representational space.
@@
We relate the p-score with performance on $\mathcal{D}_{test-OOD}$ for both the shape and color task for all models mentioned in Sections 4 and 5. Results for experiments with $8$, $64$, and $216$ colors are shown in Figure 11(a). Relating to the generalization capabilities in the shape task, we see a positive trend in o.o.d generalization when the p-score increases.
@@
To understand how significant this relationship is, we calculate the Pearson correlation between p-score and out of distribution performance on the shape task. See supplementary material for details. We obtain a 0.73 Pearson correlation with very low p-value, suggesting the strong relationship between p-score and SG.
@@
We also plot the relationship between NMI and p-score, as the positive relation both values have out-of-distribution generalization might be mediated by one or the other. Results are shown in Figure 11(b). A clear relationship between NMI and p-score appears: lower NMI datasets are related to higher p-scores in representations. Given that NMI is a stable property of the dataset, this suggests that lower NMI datasets tend to produce models with representations with higher p-scores. This suggests a novel mechanism by which lower NMI datasets produce inductive biases during learning that allow models to achieve greater out-of-distribution generalization. These results provide evidence that the dataset distribution influences the representational space geometry of the model biasing it towards out-of-distributional generalization.
(a) OOD Accuracy
(b) NMI
@@
Figure 11: (a) Out-of-distribution accuracy versus p-score for models trained on 8, 64 an 216 colors for both the shape task. A trend can be seen that increasing performance is related to higher parallellism scores in the representation. (b) NMI vs p-score for various models. Lower NMI datasets produce models with representations with higher p-score suggesting a mechanism by which lower NMI aids in achieving SG.
@@ 9 Related Work
Systematic Generalization. Many works have addressed the problem of SG, proposing a variety of strategies. These include data augmentation methods [4, 3, 64, 2], training methods [38, 11, 39], adjustments to the loss function [26], architectural changes [13, 5, 18, 36, 22, 49, 56] and neuro-symbolic approaches [44, 48]. Currently, state-of-the-art methods are based on enhanced LLMs [16, 57].
Data-driven Inductive Biases. Previous work has investigated the invariances [9], inductive biases such as shape bias [23], and robustness [20] acquired by deep learning models when trained on large-scale datasets. Prystawski et al. [55] shows the need for local structure in the training data for models to learn probabilistic reasoning abilities.
@@ 10 Conclusion
We advocate for a different way of creating inductive biases for SG on our models, by focusing on altering the training distribution over other techniques. We show three ways of manipulating the training distribution that provide significant increases in systematic generalization: diversity, burstiness and latent intervention. All of these properties enhance SG significantly, with diversity being the most important, enhancing up to an absolute 89% over the baseline, while the rest seem to be complementary to diversity.
We also study why these are effective. We show that limiting model capacity is not a factor for SG, while increased disentanglement and lower NMI between latent attributes in the dataset seem to explain most of the gains in SG. Crucially, we test for a property of brain-like representations which measures parallelism in representations under changes in a latent attribute. We find that datasets with lower NMI tend to induce more of this parallelism, a property that allows models to extrapolate by analogy. This is, to the best of our knowledge, an unreported mechanism by which dataset distribution induces SG.
@@ Acknowledgements
This work was supported by FONDECYT under Grant 1221425; and National Center for Artificial Intelligence (CENIA) under Grant FB210017 Basal ANID.
@@ 12 Correlation between P-Score and OOD Performance
In Table 1 we show the Pearson correlation between different p-scores obtained for a model and their OOD performance on the shape task. Parallelism in the representations seems to strongly correlate with better OOD performance, suggesting that parallelism may be the property that is inducing systematic generalization in the models.
@@
Table 1: Pearson Correlation between P-Scores and out-of-distribution performance on the shape task broken down per training data property. DIV: Diversity, BURST: Burstiness, LI: Latent Intervention.
@@ 13 Detailed Experimental Results
Table 2: Table displaying values to form Figure 2. Accuracy for predicting different properties for different data splits for the baseline model (8 colors). In-distribution performance for all tasks remains high, however out-of-distribution performance plummets for shape and color, suggesting the model is learning combinations of shape-color as features, instead of achieving systematic generalization. Unexpectedly, material and size also show a drop in out-of-distribution performance, even though the model has been exposed to all combinations of these attributes in training.
@@ 8 Colors
Table 3: Table displaying values to form Figure 3. Accuracy for the shape task for different amounts of training data for the baseline model (8 colors) vs model trained on 216 colors. (a) Increasing dataset size does not increase out-of-distribution performance but rather degrades it slightly. (b) With increased diversity much stronger out-of-distribution generalization is achieved, with models trained on only a quarter of the data severely outperforming the 8-color baseline.
@@
Table 4: Table displaying values to form Figure 4. Accuracy versus the number of colors in $\mathcal{D}_{train}$ for $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ for all tasks. Performance for the shape task increases drastically for the OOD split as we increase colors, increasing 86% in absolute terms over the 8-color baseline. Moreover, performance in the color task also tends to increase in the out-of-distribution split, while in-distribution only suffers slightly, even though the task becomes significantly harder. Remarkably, the material and size task rapidly increase their $\mathcal{D}_{test-OOD}$ performance as color increases.
@@
Table 6: Table displaying values to form Figure 6. Accuracy after applying latent intervention for all tasks in test-ID and test-OOD for different levels of latent intervention of the color latent attribute for various numbers of colors. Altering the color hue randomly during training allows the model to gain up to 15% more out-of-distribution accuracy over the baseline.
@@
Table 7: Table displaying values to form Figure 7. In-Distribution and Out-of-Distribution accuracy for models trained with different values for their hidden dimensions on a training set with 8 and 216 colors for all tasks.

### v3

@@ 7 Mutual Information versus Diversity
Figure 8: NMI in the training set vs OOD accuracy for the shape task for datasets with 8, 64 and 216 colors. For lower NMI, there is greater OOD performance. However, for similar levels of NMI but differing levels of diversity, we observe that greater diversity increases performance. This suggests that the effect of diversity is not only related to changing the NMI in the dataset.
Previous work [1] links Normalized Mutual Information (NMI) of latent attributes and SG. Mutual information quantifies the amount of information one variable (e.g., color) provides about another (e.g., shape). Then, this value is normalized with the entropy of the distributions to produce the NMI. The closer NMI is to zero, the more independent these factors are. In other words, a near-zero mutual information score indicates that the dataset offers strong evidence to the model that shape and color should be treated as independent attributes.
Changing diversity in a dataset impacts its NMI, so we proceed to study this effect. We seek to decouple the effect of increased diversity from a dataset’s NMI. To modulate NMI for a given level of diversity, we produce datasets that assign colors that both cubes and cylinders see during training (common colors) and a set of colors that are only seen either by cubes or by cylinders (exclusive colors). Modulating the ratio of common colors within the dataset directly alters its NMI, while keeping the number of colors constant.
To measure the NMI, we follow previous work [1] and focus on the NMI between the color and shape of all objects of the training dataset. It shows a Pearson correlation of -0.79. Figure 8 plots results for datasets with differing levels of diversity. For lower NMI, there is greater OOD performance. However, for similar levels of NMI but differing levels of diversity, we observe that greater diversity increases performance. This suggests that diversity influences more than just NMI, pointing to an additional mechanism. We hypothesize that this mechanism affects the geometry of representations, which we explore next.
@@ 8.1 Disentanglement
The role of disentaglement in promoting SG is unclear. Some studies argue that disentanglement promotes SG by enforcing factorized representations that align with the underlying structure of the data [32, 33, 18]. While other research [56, 77, 73] finds no relation between disentanglement and OOD generalization. Using DCI metrics [20], we assess measures of Disentanglement and Completeness. Disentanglement quantifies how many latent attributes each representation dimension predicts, ranging from 0 (equal predictive power across all attributes) to 1 (each dimension exclusively predicts a single attribute). Completeness measures how distributed the prediction of each attribute is across dimensions, with 0 indicating equal contribution from all dimensions and 1 meaning a single dimension predicts each attribute.
Figure 9 shows results for these metrics for representations for color and shape task, from models trained with varying color diversity. Higher diversity increases disentanglement and completeness for shape task representations, while the converse is true for color task representations.
@@
Figure 9: Disentanglement and Completeness for models trained on different levels of color diversity for vectors predicting the shape and color tasks. Both metrics increase for the shape task vectors, while the contrary happens for color task vectors.
We also complement our analysis by drawing from neuroscience-inspired studies highlighting how parallelism in representations enhances generalization [9, 35].
@@ 8.2 Parallelism in Representations
Consider the classic example for analogy in DNNs: $king-man+woman=queen$ which implies: $king-man=queen-woman=royalness$.
@@
Here, subtracting gender produces vectors that move in the same direction (an abstract property one might consider royalness) for two otherwise different vectors. Encoding this property consistently in a single direction allows the model to generalize that concept through analogy—a powerful inductive bias for solving new tasks [55].
@@
We proceed to apply this same principle to the latent factors contained in our dataset. From a starting set of 1,024 images, we sample 3,500 pairs of representations from objects that share a common latent attribute. To represent each object we use the representation vector associated to the attribute we wish to study, which in this case will be either shape or color. We proceed to find for each of these pairs another pair of representations which represents a change in this attribute (i.e. for shape, from a pair of red and green cubes, we look for a pair of red and green cylinders). We substract these pairs and from the resulting pair we obtain a single value called a p-score [35]. This score is calculated by measuring the cosine of the angle between vectors in the resulting pair. We report the mean of this value over the 3,500 pairs averaged over 5 runs.
@@
Figure 10: Computation of the p-score metric for vectors $V_{A}$ and $V_{B}$, representing the effect of ablating features on shape embeddings. We first subtract the embeddings to obtain vectors $V_{A}$ and $V_{B}$ and then calculate the degree of parallelism between them using cosine similarity. A high p-score indicates strong parallelism, suggesting that the attribute (in this case, cubeness) is encoded consistently in representational space.
@@
We relate the p-score with performance on $\mathcal{D}_{test-OOD}$ for both the shape and color task for all models mentioned in Sections 4 and 5. Results for experiments with $8$, $64$, and $216$ colors are shown in Figure 11(a). Relating to the generalization capabilities in the shape task, we see a positive trend in OOD generalization when the p-score increases.
@@
To assess how significant this relationship is, we compute the Pearson correlation between p-score and $\mathcal{D}_{test-OOD}$ performance on the shape task (details in the supplementary material). The result, $0.73$ with very low p-value, indicates a strong link between p-score and SG.
@@
We also plot the relationship between NMI and p-score, as the positive relation both values have OOD generalization might be mediated by one or the other. Results are shown in Figure 11(b). A clear relationship between NMI and p-score appears: lower NMI datasets are related to higher p-scores in representations. Given that NMI is a stable property of the dataset, this suggests that lower NMI datasets tend to produce models with representations with higher p-scores. This suggests a novel mechanism by which lower NMI datasets produce inductive biases during learning that allow models to achieve greater OOD generalization. These results provide evidence that the dataset distribution influences the representational space geometry of the model biasing it towards OOD generalization.
(a) OOD accuracy vs. p-score
(b) p-score vs. NMI
@@
Figure 11: (a) OOD accuracy on the shape task vs. p-score for models trained on 8, 64, and 216 colors. Showing a trend where higher parallelism scores in the representation correlate with better performance. (b) NMI vs p-score across models. Lower NMI datasets produce models with representations with higher p-score suggesting a mechanism by which lower NMI aids SG.
@@ 9 Related Work
Systematic Generalization.
Many works have addressed the problem of SG, proposing a variety of strategies. These include data augmentation methods [5, 3, 78, 2], training methods [44, 12, 45], adjustments to the loss function [30], architectural changes [14, 6, 19, 42, 24, 59, 67] and neuro-symbolic approaches [50, 57]. Currently, state-of-the-art methods are based on enhanced LLMs [17, 68].
Data-driven Inductive Biases.
Previous work has investigated the invariances [10], inductive biases such as shape bias [25], and robustness [22] acquired by deep learning models when trained on large-scale datasets. Prystawski et al. [66] shows the need for local structure in the training data for models to learn probabilistic reasoning abilities.
@@
Linear Representation Hypothesis
The parallelism of neural network representations is closely related to the linear representation hypothesis. This hypothesis suggests that neural network activations can be understood in terms of distinct features, which correspond to specific directions in activation space. Prior research has provided evidence for this hypothesis in both categorical [54, 58, 21, 53, 74, 4, 26, 60] and numerical features [29, 31]. Furthermore, it has been proposed that learning linear representations can enhance OOD generalization by reducing the nonlinear entanglement or contamination of different features [79].
@@ 10 Conclusion
We propose a different way of shaping inductive biases for SG, focusing on altering the properties of training data over other techniques. We show three properties of the training data that, when altered, significantly boost SG: diversity, burstiness, and latent intervention. Diversity has the strongest impact, improving SG by up to an absolute 89% over baseline, while the other factors complement it.
We also study why these are effective. We show that limiting model capacity is not a factor for SG, while increased diversity and lower NMI between latent attributes in the dataset seem to explain most of the gains in SG. Crucially, we test for a property of brain-like representations that measures parallelism in representations under changes in a latent attribute. We find that datasets with lower NMI tend to induce more of this parallelism, a property that allows models to extrapolate by analogy. This is, to the best of our knowledge, an unreported mechanism by which dataset distribution induces SG.
@@ Acknowledgements
This work was supported by FONDECYT under Grants 1221425 and 11230762; and National Center for Artificial Intelligence (CENIA) under Grant FB210017 Basal ANID. We gratefully acknowledge the support of the AMD University Program (AUP), whose computing resources enabled the experiments conducted in this work.
@@
11 Hyper Parameters
Images were resized to 224 × 224 pixels, and divided into patches of 16 × 16 pixels before being fed to the model. In the text, a masked language modeling (MLM) probability of 0.15 was used to randomly mask text tokens before feeding them to the model.
The model used during our experiments is a Transformer [75] with a hidden layer dimension of 256, with 4 transformer layers and 4 attention heads. Training was carried out using the Adam optimizer [40] with a learning rate of $1\times 10^{-4}$, a batch size of 256 and for 1000 epochs.
@@ 13 Correlation between P-Score and OOD Performance
In Table 1 we show the Pearson correlation between different p-scores obtained for a model and their OOD performance on the shape task. Parallelism in the representations seems to strongly correlate with better OOD performance, suggesting that parallelism may be the property that is inducing SG in the models.
@@
Table 1: Pearson Correlation between P-Scores and OOD performance on the shape task broken down per training data property. DIV: Diversity, BURST: Burstiness, LI: Latent Intervention.
@@ 14 Detailed Experimental Results
Table 2: Table displaying values to form Figure 2. Accuracy for predicting different properties for different data splits for the baseline model (8 colors). ID performance for all tasks remains high, however OOD performance plummets for shape and color, suggesting the model is learning combinations of shape-color as features, instead of achieving SG. Unexpectedly, material and size also show a drop in OOD performance, even though the model has been exposed to all combinations of these attributes in training.
@@ 8 Colors
Table 3: Table displaying values to form Figure 3. Accuracy for the shape task for different amounts of training data for the baseline model (8 colors) vs model trained on 216 colors. (a) Increasing dataset size does not increase OOD performance but rather degrades it slightly. (b) With increased diversity much stronger OOD generalization is achieved, with models trained on only a quarter of the data severely outperforming the 8-color baseline.
@@
Table 4: Table displaying values to form Figure 4. Accuracy versus the number of colors in $\mathcal{D}_{train}$ for $\mathcal{D}_{test-ID}$ and $\mathcal{D}_{test-OOD}$ for all tasks. Performance for the shape task increases drastically for the OOD split as we increase colors, increasing 86% in absolute terms over the 8-color baseline. Moreover, performance in the color task also tends to increase in the OOD split, while ID only suffers slightly, even though the task becomes significantly harder. Remarkably, the material and size task rapidly increase their $\mathcal{D}_{test-OOD}$ performance as color increases.
@@
Table 6: Table displaying values to form Figure 6. Accuracy after applying latent intervention for all tasks in test-ID and test-OOD for different levels of latent intervention of the color latent attribute for various numbers of colors. Altering the color hue randomly during training allows the model to gain up to 15% more OOD accuracy over the baseline.
@@
Table 7: Table displaying values to form Figure 7. ID and OOD accuracy for models trained with different values for their hidden dimensions on a training set with 8 and 216 colors for all tasks.